What is the length of a Q-switched pulse from a 10-cm long neodymium-glass laser with a 90% reflective output mirror? Assume that the refractive index of glass is 1.5.

Answers: A.

20 nanoseconds

B.

1 microsecond

C.

1 nanosecond

D.

5 nanoseconds

E.

10 nanoseconds
Answer #1

Result : (E) 10 nanoseconds

Explanation :

  • Given : L = 10 cm ; n = 1.5 ; R = 90% (0.90)

The length of a Q switched pulse is given by :

\frac {2Ln}{c(1-R)}

= \frac {2(10*10^{-2})(1.5)}{(3*10^8)(1-0.90)}

= 10× 10-9 s

= 10 ns